OpenStudy (anonymous):

How many grams are in 6.7 x 10 18 formula units of sodium acetate?Show all your work for credit and label all your numbers.

OpenStudy (anonymous):

is it 6.7 x 10^18?

OpenStudy (matt101):

A formula unit is the "molecule" you have from the empirical formula of a compound. In the case of sodium acetate, the formula unit is NaCH3COO. You have (6.7 *10^18)/(6.02 * 10^23) = 1.11 * 10^(-5) moles of NaCH3COO. The molar mass of NaCH3COO is 82 g/mol, so you have 82 * 1.11 * 10^(-5) = 9.1 * 10^(-4) g, or 0.91 mg. If you need me to clarify something let me know.

OpenStudy (anonymous):

@Tttaammmaannnaa yes!